# MagR: Weight Magnitude Reduction for Enhancing Post-Training Quantization

### v1

@@ Abstract
In this paper, we present a simple optimization-based preprocessing technique called Weight Magnitude Reduction (MagR) to improve the performance of post-training quantization. For each linear layer, we adjust the pre-trained floating-point weights by solving an $\ell_{\infty}$-regularized optimization problem. This process greatly diminishes the maximum magnitude of the weights and smooths out outliers, while preserving the layer’s output. The preprocessed weights are centered more towards zero, which facilitates the subsequent quantization process. To implement MagR, we address the $\ell_{\infty}$-regularization by employing an efficient proximal gradient descent algorithm. Unlike existing preprocessing methods that involve linear transformations and subsequent post-processing steps, which can introduce significant overhead at inference time, MagR functions as a non-linear transformation, eliminating the need for any additional post-processing. This ensures that MagR introduces no overhead whatsoever during inference. Our experiments demonstrate that MagR achieves state-of-the-art performance on the Llama family of models. For example, we achieve a Wikitext2 perplexity of 5.95 on the LLaMA2-70B model for per-channel INT2 weight quantization without incurring any inference overhead.
@@ 1 Introduction
The enormous computational demands for pre-training and fine-tuning Large Language Models (LLMs) have led to the emergence of Post-Training Quantization (PTQ) behdin2023quantease ; frantar2022optimal ; li2021brecq ; lin2023bit ; maly2023simple ; nagel2019data ; wang2022deep ; wei2022qdrop ; zhang2024comq ; zhang2023post as a promising solution for quantizing these models. Unlike Quantization Aware Training (QAT) cai2017deep ; choi2018pact ; hubara2018quantized ; rastegari2016xnor ; yin2019blended , which is designed to minimize a global training loss for quantization parameters, PTQ directly applies low-precision calibration to a pre-trained full-precision model using a minimal set of calibration samples. By aiming to identify an optimal quantized model locally through the minimization of a simplified surrogate loss, PTQ offers computational savings and resource efficiency compared to QAT. However, PTQ often lags behind QAT in accuracy, particularly for ultra-low precision lower than 4 bit. Thus, it remains an open problem to achieve an improved balance between cost and performance for PTQ-based approaches.
Motivation. To achieve state-of-the-art performance, the latest advances in PTQ chee2024quip ; lin2023awq ; ma2024affinequant ; shao2023omniquant ; wei2023outlier have proposed applying a linear transformation to process the pre-trained weights within a linear layer. This strategy of linear transformation aims to make the weights more suitable for the subsequent quantization procedure by reducing their magnitudes and suppressing outliers. In a nutshell, given the features $\boldsymbol{X}$ and weights $\boldsymbol{W}$, one constructs some matrix $\boldsymbol{T}$ such that $\boldsymbol{T}\boldsymbol{W}$ is better conditioned than $\boldsymbol{W}$ in terms of being quantization-friendly. Such designs of $\boldsymbol{T}$ include diagonal matrix (so-called channel-wise scaling) shao2023omniquant ; wei2023outlier as in AWQ lin2023awq and random orthogonal matrix as in QuIP chee2024quip . Then, quantization is performed on $\boldsymbol{T}\boldsymbol{W}$ instead of the original weights $\boldsymbol{W}$. To preserve the layer’s output, however, the inverse transformation $\boldsymbol{T}^{-1}$ has to be in turn applied to the features $\boldsymbol{X}$, namely,
@@
To address this problem, we propose a simple optimization-based technique called Weight Magnitude Reduction (MagR). MagR functions as a non-linear transformation on weights without altering the original layer-wise features. The optimization program is designed to find new weights with minimal maximum magnitude, i.e., the $\ell_{\infty}$ norm, while preserving the layer’s outputs.
Contributions. We propose a non-linear approach, MagR, based on $\ell_{\infty}$-regularization, to reduce the quantization scale, facilitating subsequent weight quantization while requiring no post-processing or inference overhead. See Figure 1 for comparing weight magnitudes before and after applying MagR. To efficiently solve the resulting $\ell_{\infty}$-regularization problem, we develop a proximal gradient descent algorithm which involves computing $\ell_{1}$-ball projection during the iterations. Our results on INT weight-quantization demonstrate that MagR can significantly boost accuracy in the sub-4bit regime when combined with fast gradient-free methods for layer-wise PTQ, such as rounding-to-nearest (RTN) nagel2020up and OPTQ frantar2022optq . This approach achieves performance for weight quantization comparable to state-of-the-art PTQ methods on natural language processing (NLP) tasks, including gradient-based methods using block-wise reconstruction.
@@ 2 Related Work
Recently, as the sizes of language models are exploding, there has been growing interest in developing post-training quantization (PTQ) methods chee2024quip ; frantar2022optq ; lin2023awq ; ma2024affinequant ; shao2023omniquant ; yao2024exploring ; xiao2023smoothquant for large-scale AI models like large language models (LLMs) to reduce the model sizes and accelerate inference by representing weight matrices in low precision. PTQ methods directly find the low-precision representation of the model without re-training, thereby preferred by extreme large-scale AI models. The OPTQ frantar2022optq uses approximate second-order information to calibrate the quantization. The method successfully compresses LLMs into 3 or 4 bits and can achieve reasonable accuracy in 2 bits. Researchers have found that the extreme values and the distribution of the weight entries highly affect the quantization errors and the quantized model quality. The original weight can be converted into a more quantization-friendly one by linear transformations. The approach can significantly reduce the quantization errors while bringing more time overhead during inference because of the linear transformation. OmniQuant shao2023omniquant proposes learnable weight clippings and equivalent transformations to avoid the influence of extreme values. AWQ lin2023awq searches for the most significant entries in the weight by looking at the activation and selects the scales that protect these entries. SmoothQuant xiao2023smoothquant passes the difficulty in activation quantization to weights by an equivalent linear transformation. QuIP chee2024quip and AffineQuant ma2024affinequant apply a linear transformation before quantization to make the transformed weight quantization-friendly. These approaches achieve high performance for extreme bits, like 2 bits, but introduce additional inference overhead though the transformation is carefully designed to be efficient. OmniQuant shao2023omniquant and AffineQuant ma2024affinequant can be adopted for weight-activation quantization by considering the activations in the proposed methods. The work yao2024exploring introduces a low-rank compensation method on top of other quantization methods, which employs low-rank matrices to reduce quantization errors with a minimal increase in model size.
@@ 3 Background
Notations. We denote vectors by bold small letters and matrices by bold capital ones. For a positive integer $n$, $[n]:=\{1,2,\dots,n\}$ denotes the set containing all positive integers up to $n$. For any two vectors $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{n}$, $\langle\boldsymbol{x},\boldsymbol{y}\rangle:=\sum_{i=1}^{n}x_{i}y_{i}$ is the inner product. We denote by $\|\boldsymbol{x}\|:=\sqrt{\langle\boldsymbol{x},\boldsymbol{x}\rangle}=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}$ the Euclidean norm; $\|\boldsymbol{x}\|_{1}:=\sum_{i=1}^{n}|x_{i}|$ is the $\ell_{1}$-norm; $\|\boldsymbol{x}\|_{\infty}:=\max_{1\leq i\leq n}|x_{i}|$ is the $\ell_{\infty}$-norm. For any matrix $\boldsymbol{X}\in\mathbb{R}^{m\times n}$, $\boldsymbol{X}^{\top}\in\mathbb{R}^{n\times m}$ is the transpose. Its Frobenius norm is given by $\|\boldsymbol{X}\|_{\mathrm{F}}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}X_{i,j}^{2}}$. Moreover, for vectors $\boldsymbol{x}$ and $\boldsymbol{y}$, $\boldsymbol{x}\odot\boldsymbol{y}:=(x_{1}y_{1},\dots,x_{n}y_{n})\in\mathbb{R}^{n}$ denotes the Hadamard or element-wise product, and likewise for two matrices.
@@
The most straightforward PTQ technique, known as RTN, involves directly rounding the weight matrix $\boldsymbol{W}$ without utilizing any additional data. An improvement over RTN was introduced by AWQ lin2023awq , which enhances the quantization process by incorporating channel-wise scaling on $\boldsymbol{W}$. Thanks to the simplicity of the layer-wise formulation (1), several efficient gradient-free algorithms behdin2023quantease ; frantar2022optq ; zhang2024comq ; zhang2023post have been recently proposed to address layer-wise quantization, including OPTQ. Built on top of OPTQ, QuIP subjects $\boldsymbol{X}$ and $\boldsymbol{W}$ to random orthogonal transformations to produce “incoherent" weight and Hessian matrices, leading to superior accuracy with sub-4bit quantization. However, this advantage comes with a trade-off; during inference, QuIP requires random orthogonal transformations on the feature inputs of linear layers, rendering noticeably slower throughput compared to OPTQ.
@@ 4 The Proposed Method
In this section, we present the Weight Magnitude Reduction (MagR) method based on $\ell_{\infty}$-norm regularization. which is performed immediately prior to the quantization step within each linear layer.
4.1 Rank-Deficient Feature Matrix
@@
To illustrate the idea behind the proposed MagR method, let us consider a pre-trained weight vector $\hat{\boldsymbol{w}}\in\mathbb{R}^{m}$ of a linear layer and the associated feature input matrix $\boldsymbol{X}$. MagR leverages the fact that the feature matrix $\boldsymbol{X}$ across all layers of LLMs is approximately rank-deficient. Specifically, if $\boldsymbol{X}$ is rank-deficient, the linear system modeling the layer’s output, $\boldsymbol{X}\boldsymbol{w}=\boldsymbol{X}\hat{\boldsymbol{w}}$ with variables $\boldsymbol{w}$, generally has infinitely many solutions. That is, for any $\boldsymbol{\nu}$ in the non-trivial kernel space of $\boldsymbol{X}$, we have that $\boldsymbol{w}=\hat{\boldsymbol{w}}+\boldsymbol{\nu}$ preserves the layer’s output. Among all solutions, MagR aims to identify the weight vector $\boldsymbol{w}$ with the smallest extreme value in magnitude.
@@
In chee2024quip , the authors empirically observed that the Hessian matrix $\boldsymbol{H}=\boldsymbol{X}^{\top}\boldsymbol{X}$ is approximately low-rank across all layers in open pre-trained (OPT) models zhang2022opt . Specifically, it was shown that the fraction of eigenvalues of $\boldsymbol{H}$ that satisfy $\lambda(\boldsymbol{H})>0.01\cdot\lambda_{\max}(\boldsymbol{H})$, i.e. approximate fraction rank, is much smaller than 1. Here we further examined the feature matrix of LLaMA models touvron2023llama ; touvron2023llama2 . Our approximate fraction rank of the feature matrix $\boldsymbol{X}$ is defined as the fraction of singular values of $\boldsymbol{X}$ such that $\sigma(\boldsymbol{X})>0.01\cdot\sigma_{\max}(\boldsymbol{X})$. It is worth noting that our definition of being rank-deficient is more stringent compared to the one used in chee2024quip . Indeed, if a singular value of $\boldsymbol{X}$ obeys $\sigma(\boldsymbol{X})\leq 0.01\cdot\sigma_{\max}(\boldsymbol{X})$, then the corresponding eigenvalue of $\boldsymbol{H}$ gives $\lambda(\boldsymbol{H})\leq 0.01^{2}\cdot\lambda_{\max}(\boldsymbol{H})$. Table 1 illustrates that all feature matrices extracted from LLaMA models are rank-deficient according to this definition.
@@
Table 1: The statistics of (approximate) fraction ranks in percentage (%) of feature matrix $\boldsymbol{X}$ across all layers of LLaMA models. All feature matrices are approximately rank-deficient with a fraction rank strictly less than 1 (or 100%). Some of them are highly low-rank with a fraction rank $\approx 1\%$, while some others are barely rank-deficient.

### v2

@@ Abstract
In this paper, we present a simple optimization-based preprocessing technique called Weight Magnitude Reduction (MagR) to improve the performance of post-training quantization. For each linear layer, we adjust the pre-trained floating-point weights by solving a channel-wise $\ell_{\infty}$-regularized optimization problem. This process greatly diminishes the maximum magnitude of the weights and smooths out outliers, while preserving the layer’s output. The preprocessed weights exhibit reduced range, which facilitates the subsequent quantization process. To implement MagR, we address the $\ell_{\infty}$-regularization by employing an efficient proximal gradient descent algorithm. Unlike existing preprocessing methods that involve linear transformations and subsequent post-processing steps, which can introduce significant overhead at inference time, MagR functions as a non-linear transformation, eliminating the need for any additional post-processing. This ensures that MagR introduces no overhead whatsoever during inference. Our experiments demonstrate that MagR achieves state-of-the-art performance on the Llama family of models. For example, we achieve a Wikitext2 perplexity of 5.95 on the LLaMA2-70B model for per-channel INT2 weight quantization without incurring any inference overhead. The code is available at https://github.com/AozhongZhang/MagR
@@ 1 Introduction
The enormous computational demands for pre-training and fine-tuning Large Language Models (LLMs) have led to the emergence of Post-Training Quantization (PTQ) behdin2023quantease ; frantar2022optimal ; li2021brecq ; lin2023bit ; maly2023simple ; nagel2019data ; wang2022deep ; zhang2024comq ; zhang2023spfq ; zhang2023post ; wu2024ptq4dit ; wang2024quest , as a promising solution for quantizing these models. Unlike Quantization Aware Training (QAT) cai2017deep ; choi2018pact ; courbariaux2015binaryconnect ; hubara2018quantized ; li2016ternary ; li2023feature ; yin2019understanding ; yin2018binaryrelax ; yin2019blended ; yin2016quantization , which is designed to minimize a global training loss for quantization parameters, PTQ directly applies low-precision calibration to a pre-trained full-precision model using a minimal set of calibration samples. By aiming to identify an optimal quantized model locally through the minimization of a simplified surrogate loss, PTQ offers computational savings and resource efficiency compared to QAT. However, PTQ often lags behind QAT in accuracy, particularly for ultra-low precision lower than 4-bit. Thus, it remains an open problem to achieve an improved balance between cost and performance for PTQ-based approaches.
Motivation. To achieve state-of-the-art performance, the latest advances in PTQ chee2024quip ; lin2023awq ; ma2024affinequant ; shao2023omniquant ; wei2023outlier have proposed applying a linear transformation to process the pre-trained weights within a linear layer. This strategy of linear transformation aims to make the weights more suitable for the subsequent quantization procedure by reducing their magnitudes and suppressing outliers. In a nutshell, given the features $\boldsymbol{X}$ and weights $\boldsymbol{W}$, one constructs linear transformation $\boldsymbol{T}$ such that $\boldsymbol{T}\boldsymbol{W}$ is better conditioned than $\boldsymbol{W}$ in terms of being quantization-friendly. Such designs of $\boldsymbol{T}$ include diagonal matrices (so-called channel-wise scaling) lin2023awq ; shao2023omniquant ; wei2023outlier , random transformations chee2024quip ; tseng2024quip , and finite frames adepu2024framequant ; czaja2024frame . Then, quantization is performed on $\boldsymbol{T}\boldsymbol{W}$ instead of the original weights $\boldsymbol{W}$. To preserve the layer’s output, however, the inverse transformation $\boldsymbol{T}^{-1}$ has to be in turn applied to the features $\boldsymbol{X}$, namely,
@@
To address this problem, we propose a simple optimization-based technique called Weight Magnitude Reduction (MagR). MagR functions as a non-linear transformation on weights without altering the original features/activations. The optimization program is designed to find new weights with minimal maximum magnitude, i.e., the $\ell_{\infty}$ norm, while preserving the layer’s outputs.
Contributions. We propose a non-linear approach, MagR, based on channel-wise $\ell_{\infty}$-regularized least squares, to reduce the quantization scale without compromising the performance of pre-trained model, facilitating subsequent weight quantization while requiring no post-processing or inference overhead. See Figure 1 for comparing weight magnitudes before and after applying MagR. To address the $\ell_{\infty}$-regularization problem, we develop an efficient and parallelizable proximal gradient descent algorithm that involves computing $\ell_{1}$-ball projections at each iteration. Specifically, MagR preprocessing on a single Nvidia A100 GPU takes merely 15 min for LLaMA2-7B and 3.5 hr for the 70B model. Our results on INT weight-quantization demonstrate that MagR can significantly boost the performance in the sub-4bit regime when combined with fast gradient-free methods for layer-wise PTQ, such as rounding-to-nearest (RTN) nagel2020up and OPTQ frantar2022optq . This approach achieves performance for weight quantization at least comparable to state-of-the-art PTQ methods on natural language processing (NLP) tasks, including gradient-based methods using block-wise reconstruction.
@@ 2 Related Work
Recently, as the sizes of language models are exploding, there has been growing interest in developing post-training quantization (PTQ) methods chee2024quip ; frantar2022optq ; lin2023awq ; ma2024affinequant ; shao2023omniquant ; xiao2023smoothquant ; yao2024exploring for large-scale AI models like large language models (LLMs) to reduce the model sizes and accelerate inference by representing weight matrices in low precision. PTQ methods directly find the low-precision representation of the model without re-training, thereby preferred by extreme large-scale AI models. The OPTQ frantar2022optq uses approximate second-order information to calibrate the quantization. The method successfully compresses LLMs into 3 or 4 bits and can achieve reasonable accuracy in 2 bits. Researchers have found that the extreme values and the distribution of the weight entries highly affect the quantization errors and the quantized model quality. The original weight can be converted into a more quantization-friendly one by linear transformations. The approach can significantly reduce the quantization errors while bringing more time overhead during inference because of the linear transformation. OmniQuant shao2023omniquant proposes learnable weight clippings and equivalent transformations to avoid the influence of extreme values. AWQ lin2023awq searches for the most significant entries in the weight by looking at the activation and selects the scales that protect these entries. SmoothQuant xiao2023smoothquant passes the difficulty in activation quantization to weights by an equivalent linear transformation. QuIP chee2024quip , AffineQuant ma2024affinequant and FrameQuant adepu2024framequant apply a linear transformation before quantization to make the transformed weight quantization-friendly. These approaches achieve high performance for extreme bits, like 2 bits, but introduce additional inference overhead though the transformation is carefully designed to be efficient. OmniQuant shao2023omniquant and AffineQuant ma2024affinequant can be adopted for weight-activation quantization by considering the activations in the proposed methods. The work yao2024exploring introduces a low-rank compensation method on top of other quantization methods, which employs low-rank matrices to reduce quantization errors with a minimal increase in model size. By modeling the quantization residual as an $\ell_{\infty}$-bounded perturbation, alizadeh2020gradient proposes applying an $\ell_{1}$ penalty on the gradient of loss to enhance quantization robustness.
The works most closely related to ours are kundu2023r2 and maly2023simple , both utilizing $\ell_{\infty}$ norm to regularize or constrain the weight range to a smaller scale. The Range Regularization (R2) method kundu2023r2 applies an $\ell_{\infty}$ penalty or its variants to the conventional network loss to regularize the weight range during end-to-end model pre-training, optimized via SGD. However, this approach becomes practically infeasible for large-scale models. In maly2023simple , a layer-wise pre-processing technique is proposed, which involves solving an intractable $\ell_{0}$-minimization problem while constraining the $\ell_{\infty}$-norm of weights.
@@ 3 Background
Notations. We denote vectors by bold small letters and matrices by bold capital ones. For a positive integer $n$, $[n]:=\{1,2,\dots,n\}$ denotes the set containing all positive integers up to $n$. For any two vectors $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{n}$, $\langle\boldsymbol{x},\boldsymbol{y}\rangle:=\sum_{i=1}^{n}x_{i}y_{i}$ is the inner product. We denote by $\|\boldsymbol{x}\|:=\sqrt{\langle\boldsymbol{x},\boldsymbol{x}\rangle}=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}$ the Euclidean norm; $\|\boldsymbol{x}\|_{1}:=\sum_{i=1}^{n}|x_{i}|$ is the $\ell_{1}$-norm; $\|\boldsymbol{x}\|_{\infty}:=\max_{1\leq i\leq n}|x_{i}|$ is the $\ell_{\infty}$-norm. For any matrix $\boldsymbol{X}\in\mathbb{R}^{m\times n}$, $\boldsymbol{X}^{\top}\in\mathbb{R}^{n\times m}$ is the transpose. We denote the spectrum norm of $\boldsymbol{X}$ by $\|\boldsymbol{X}\|=\sigma_{\max}(\boldsymbol{X})$, which equals its maximum singular value. Its Frobenius norm is given by $\|\boldsymbol{X}\|_{\mathrm{F}}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}X_{i,j}^{2}}$. Moreover, for vectors $\boldsymbol{x}$ and $\boldsymbol{y}$, $\boldsymbol{x}\odot\boldsymbol{y}:=(x_{1}y_{1},\dots,x_{n}y_{n})\in\mathbb{R}^{n}$ denotes the Hadamard or element-wise product, and likewise for two matrices.
@@
The most straightforward PTQ technique, known as RTN, involves directly rounding the weight matrix $\boldsymbol{W}$ without utilizing any additional data. An improvement over RTN was introduced by AWQ lin2023awq , which enhances the quantization process by incorporating channel-wise scaling on $\boldsymbol{W}$.
Thanks to the simplicity of the layer-wise formulation (1), several efficient gradient-free algorithms behdin2023quantease ; frantar2022optq ; zhang2024comq ; zhang2023post have been recently proposed to address layer-wise quantization, including OPTQ. Built on top of OPTQ, QuIP subjects $\boldsymbol{X}$ and $\boldsymbol{W}$ to random orthogonal transformations to produce “incoherent" weight and Hessian matrices, leading to superior accuracy with sub-4bit quantization. However, this advantage comes with a trade-off; during inference, QuIP requires random orthogonal transformations on the feature inputs of linear layers, rendering noticeably slower throughput compared to OPTQ.
@@ 4 The Proposed Method
In this section, we present the Weight Magnitude Reduction (MagR) method based on $\ell_{\infty}$-norm regularization, which is applied just before the quantization step within each linear layer. The intuition behind MagR is based on the following simple estimate of the layer-wise quantization error. Given the feature/activation matrix $\boldsymbol{X}$, the quantizer $\mathcal{Q}$, and any pre-trained weights $\boldsymbol{w}\in\mathbb{R}^{m}$, we have:
$\min_{\boldsymbol{w}_{q}\in\mathbb{Q}}\|\boldsymbol{X}\boldsymbol{w}_{q}-\boldsymbol{X}\boldsymbol{w}\|\leq\|\boldsymbol{X}(\mathcal{Q}(\boldsymbol{w})-\boldsymbol{w})\|\leq\|\boldsymbol{X}\|\|\mathcal{Q}(\boldsymbol{w})-\boldsymbol{w}\|\leq\frac{\sigma_{\max}(\boldsymbol{X})\sqrt{m}}{2}\delta,$
where $\delta=\frac{\max(\boldsymbol{w})-\min(\boldsymbol{w})}{2^{b}-1}$ is the quantization step size. This shows that reducing the range of weights helps to suppress the quantization error. With this in mind, MagR preprocessing is designed to achieve two key effects:
  • First, it effectively reduces the channel-wise (or column-wise) maximum magnitude of the weights, as illustrated by Figure 1.
  • Second, it preserves the model’s original performance with minimal accuracy loss after preprocessing. Table 1 demonstrates that MagR preprocessing maintains the perplexity of the pre-trained models, with only minor degradation.
Table 1: A comparison of perplexity (PPL) for the original pre-trained and the MagR-processed LLaMA2 models.
[…]
LLaMA2-7B Original 5.47 6.97
MagR 5.52 7.04
LLaMA2-13B Original 4.88 6.46
MagR 4.92 6.52
LLaMA2-70B Original 3.31 5.52
MagR 3.35 5.56
4.1 Approximately Rank-Deficient Feature Matrix
@@
To illustrate the idea behind the proposed MagR method, let us consider a pre-trained weight vector $\hat{\boldsymbol{w}}\in\mathbb{R}^{m}$ of a linear layer and the associated feature input matrix $\boldsymbol{X}$. MagR leverages the fact that the feature matrix $\boldsymbol{X}$ across all layers of LLMs is approximately rank-deficient. Specifically, if $\boldsymbol{X}$ is exactly rank-deficient, the linear system modeling the layer’s output, $\boldsymbol{X}\boldsymbol{w}=\boldsymbol{X}\hat{\boldsymbol{w}}$ with variables $\boldsymbol{w}$, generally has infinitely many solutions. That is, for any $\boldsymbol{\nu}$ in the non-trivial kernel space of $\boldsymbol{X}$, we have that $\boldsymbol{w}=\hat{\boldsymbol{w}}+\boldsymbol{\nu}$ preserves the layer’s output. Among all solutions, MagR aims to identify the weight vector $\boldsymbol{w}$ with the smallest extreme value in magnitude.
@@
In chee2024quip , the authors empirically observed that the Hessian matrix $\boldsymbol{X}^{\top}\boldsymbol{X}$ is approximately low-rank across all layers in open pre-trained (OPT) models zhang2022opt . Here we examined the feature matrix of LLaMA models touvron2023llama ; touvron2023llama2 . Our approximate fraction rank of the feature matrix $\boldsymbol{X}$ is defined as the fraction of singular values of $\boldsymbol{X}$ such that $\sigma(\boldsymbol{X})>0.01\cdot\sigma_{\max}(\boldsymbol{X})$. Table 2 illustrates that all feature matrices extracted from LLaMA models are indeed rank-deficient according to this definition.
@@
Table 2: The statistics of (approximate) fraction ranks in percentage (%) of feature matrix $\boldsymbol{X}$ across all layers of LLaMA models. All feature matrices are approximately rank-deficient with a fraction rank less than 100%. Some of them are highly low-rank with a fraction rank $\approx 1\%$.